A stone is dropped into a well and 4 seconds later the sound of the splash is heard. If sounds travel at a speed of 340 m/s, how deep is the well ?


1
Expert's answer
2022-02-13T12:16:15-0500

Given:

"t=4\:\rm s"

"g=9.8\:\rm m/s^2"

"u=340\:\rm m/s"


The time between you dropped a stone and you heard the sound of splash


"t=t_1+t_2=\sqrt{2h/g}+h/u""4=\sqrt{2h/9.8}+h/340"

The root of this equation


"h=70.5\:\rm m"